OpenStudy (anonymous):

It is given that x varies directly as y and inversely as z^(2). If y is decreased by 10% and z is increased by 20%, then x is decreased by __%?

OpenStudy (anonymous):

From it equation can be formed as : \[x = \frac{ky}{z^2}\]

OpenStudy (anonymous):

x=0.9yk / 1.44z^(2) ?

OpenStudy (anonymous):

If y is decreased by 10 % this means: \[y' = y - \frac{10}{100}y = \frac{90}{100}y \implies \frac{9}{10}y\] Similarly: \[z' = \frac{120}{100}z \implies \frac{6}{5}z\]

OpenStudy (anonymous):

Now x has become : \[\large x = \frac{k \times \frac{9}{10}y}{\frac{36}{25}z^2}\]

OpenStudy (anonymous):

*x'

OpenStudy (anonymous):

Now just divide both x and x'

OpenStudy (anonymous):

(x' - x )/x * 100%

OpenStudy (anonymous):

i am problem in calculating ...

OpenStudy (anonymous):

You are left with : \[\frac{x}{x'} = \frac{0.9}{1.44 \times 1.44} = 0.434\]

OpenStudy (anonymous):

what??

OpenStudy (anonymous):

Just reverse it.. \[\frac{x'}{x} = 2.304\]

OpenStudy (anonymous):

Electricity gone, have to log out sorry..

OpenStudy (anonymous):

never mind, thanks

OpenStudy (goformit100):

@kryton1212 Well Done :)

OpenStudy (anonymous):

i haven't done anything..

OpenStudy (anonymous):

\[\frac{ x*z^2 }{ y }=k\] k is a constant.. if y decrease %10 then x dectease %10, i z increase %20, ( it means \[z^{2}\] incresse %4, and it means x decrease %4 more.. but this %4 is the %90's %4.. in brief %4*%90= %3.6.. ans substract this value from %90, you get %86.4.. this is the last value of x.. it means x decteased (%100-86.4) %13.6 totally..

Directrix (directrix):

@kryton1212 If you see this, post the question here.

Directrix (directrix):

>>hey, the openstudy page have loaded a very long time, can i ask you some questions about Maths? @kryton1212 The OpenStudy site has been down for 8 or more hours. No word given on Facebook on the OS page as to when to expect functionality.